40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4904     Entry time: Tue Jun 28 22:36:04 2011
Author: Jamie 
Type: Update 
Category: SUS 
Subject: Checking binary switching of SUS whitening filter 

I have been checking the binary output switching for the SUS whitening filters. It appears that the whitening switching is working for (almost) all the vertex suspensions (BS, ITMX, ITMY, PRM, SRM), but not for the ETMs.

The table below lists the output from my switch-checking script (attached). The script uses the SUS digital lockin to drive one coil and measure the same coil's OSEM response, repeating for each coil/OSEM pair. I used a lockin drive frequency of about 10 Hz, at which the whitening filter should have 10 db of gain.

All but one of the vertex OSEMS show the proper response (~10db gain at 10Hz) when the whitening is switched on from the digital controls. ITMY UL appears to not be switching, which I fear is due to my electronics fail noted in my previous log post.  The ETMs are clearly not switching at all.

I will try to get the ETM switching working tomorrow, as well as try to asses what can be done about the ITMY UL switch.  After that I will work on confirming the coil drive dewhite switching.

lockin settings

freq: 10.123 Hz
amp: 10000
I/Q filters: 0.1 Hz LP, 4-pole butterworth

response

BS
ul : 3.31084503062 = 10.3987770676 db
ll : 3.34162124753 = 10.4791444741 db
sd : 3.43226254574 = 10.7116100229 db
lr : 3.28602651913 = 10.3334212798 db
ur : 3.29361593249 = 10.3534590969 db

ITMX
ul : 3.37499773336 = 10.5654697099 db
ll : 3.2760924572  = 10.3071229966 db
sd : 3.13374799272 =  9.9212813757 db
lr : 3.28133776018 = 10.3210187243 db
ur : 3.37250879937 = 10.5590618297 db

ITMY
ul : 0.99486434364 = -0.0447226830807 db
ll : 3.39420873724 = 10.6147709414 db
sd : 3.88698713176 = 11.7922620572 db
lr : 3.357123865   = 10.5193473069 db
ur : 3.37876008179 = 10.5751470918 db

PRM
ul : 3.26758918055 = 10.2845489876 db
ll : 3.32023820566 = 10.4233848529 db
sd : 3.25205538857 = 10.2431586766 db
lr : 3.24610681962 = 10.227256141  db
ur : 3.31311970305 = 10.4047425446 db

SRM
ul : 3.30506423619 = 10.3835980943 db
ll : 3.28152094133 = 10.3215036019 db
sd : 3.08566647696 =  9.7869796462 db
lr : 3.30298270419 = 10.378125991  db
ur : 3.3012249406  = 10.3735023505 db

ETMX
ul : 0.99903400106 = -0.00839461539757 db
ll : 0.99849991349 = -0.0130393683795 db
sd : 1.00314092883 =  0.0272390056874 db
lr : 1.00046493718 =  0.00403745453682 db
ur : 1.00265600785 =  0.0230392084558 db

ETMY
ul : 1.00223179107 =  0.0193634913327 db
ll : 0.96755532811 = -0.286483823189 db
sd : 1.00861855271 =  0.0745390477589 db
lr : 1.05718545676 =  0.483023602007 db
ur : 0.99777406174 = -0.0193558045143 db
Attachment 1: botest.py  5 kB  | Hide | Hide all
#!/usr/bin/env python

import sys
import os
import subprocess
import time
import pickle
from numpy import *
import nds
import matplotlib
... 207 more lines ...
ELOG V3.1.3-